# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I have digitized a unique daily longitudinal survey dataset on household budgets to analyze how consumption is impacted by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
Keywords: consumption smoothing; credit purchases; daily budget data; local retailers; measurement error; risk-coping strategy; risk-sharing
@@ 1 Introduction
Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). A series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), Ravallion and Chaudhuri (1997), Grimard (1997), and Fafchamps and Lund (2003). Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. The households are vulnerable to risk, and the idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets. Therefore, consumption smoothing is an essential agenda for mitigating poverty, especially in developing economies. [^2] Smoothness of consumption is also an important factor in human capital formation. For example, Foster (1995) shows that children’s growth patterns are influenced by the inefficiency of the credit markets using the floods of 1988 in Bangladesh as a shock. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall shocks in early life increased the survival probabilities of the girls in rural India. Gertler and Gruber (2002) also offers evidence that the smoothing of consumption against health shocks is fairly imperfect in Indonesia.
Historical economies offer experimental environments to investigate the means and degrees of consumption smoothing under less developed societies. Lessons from historical institutions in developed economies shall provide meaningful insights for today’s developing economies. Despite this advantage, little has been known about historical consumption smoothing behaviors, partially because of the lack of longitudinal household budget datasets. This study digitized a unique daily longitudinal budget survey on working-class households conducted in Tokyo around 1919 and employed systematic econometric methods to explore the consumption behaviors among factory worker households in a representative manufacturing area. The result shows that the full insurance hypothesis is strongly rejected for overall consumption expenditure. The estimated income elasticity is close to that for the rural households in southern India around 1980. This means that the factory worker households in historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, I also found evidence that while expenses on dispensable categories are likely to be cut in response to the idiosyncratic shocks, expenses on indispensable consumption categories are partially smoothed in the short-run. Mechanism analyses provide suggestive evidence that credit purchases in the local economy played an important role in mitigating adverse idiosyncratic shocks.
@@
The first contribution of this study is to offer a case study of a metropolis in a developing economy. Urbanization is now a general phenomenon in the world, and the shares of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This fact raises the need for a deeper understanding of consumption behavior among urban working households in the process of economic development. Despite this, there is still little evidence for the urban worker households in the literature. While Townsend (1995) offers the results for Bangkok in Thailand between 1975 and 1990, a large part of the existing studies focused on the rural economies (Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). [^3] Recent field experimental studies also focus on the village economies (Comola and Prina 2023; Macours et al. 2022). A few exceptional studies provide cases on the middle-sized contemporary economies (Skoufias and Quisumbing 2005; Aguila et al. 2017). In historical context, Ogasawara (2024) analyzes the factory workers’ households in Osaka City around 1920, which was roughly half the size of Tokyo City (Online Appendix Table A.1). This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing the risk-sharing behaviors in the historical metropolis. [^4] Focusing on the sample sharing the similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). I found evidence of consumption smoothing using credit transactions with local retailers among working-class households.
Macroeconomics literature has investigated the theoretical mechanisms behind the false rejections of the null of the risk-sharing hypothesis. An essential factor disturbing the valid statistical inference is measurement error in aggregate data (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). This study reveals a specific generating process of systematic measurement errors in the aggregated income and expenditure series in the survey dataset. In my empirical setting, the factory workers had two paydays in a month: one was set in the middle and the other at the end of the month. Hence, the panel dataset based on the calendar months leads to a substantial mismatch between the aggregate income and expenditure series. [^5] If payday is set soon after the initial date in a certain time cell, it does not cause a mismatch in income and expenditure. However, when payday is set just before the initial date in the cell, it leads to measurement errors because the consumption schedule in that cell shall based on the earnings in the previous cell. In other words, this way of matching assumes that the households in a certain cell make their consumption schedule based on future earnings, which is unrealistic. I found that the aggregation ignoring this mechanism causes false rejections in most consumption subcategories due to the attenuation bias. This evidence shows the importance of reflecting the timing of paydays in the aggregation and invokes careful reviews of the results from the aggregate budget datasets.
@@
The second contribution is to add a new analytical perspective to economic history studies. How working-class households overcame their economic hardships during the late 19th to the early 20th century Britain has attracted wider attention. Weights of evidence from these studies have revealed the roles of historical financial institutions and labor supply adjustments in compensating income losses in working-class households. [^6] Horrell and Oxley (2000) found the important roles of the labor supply adjustments among family members in coping with the income losses of the head in working-class households in late 19th-century Britain. Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractural savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use the savings institutions. Kiesling (1996) and Boyer (1997) focused on the risk-coping behaviors among the cotton workers households in Lancashire during the recession due to Civil War. Another strand of literature in financial history provided supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005). Despite this, there is little evidence of how those risk-coping strategies contributed to smooth household consumption. In addition, the previous studies have utilized the cross-sectional variations in income to estimate the elasticities for the temporary income sources. [^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings to sort out the shocks. While Scott and Walker (2012) considers the smoothness of expenditure rather than income based on the deviation from the mean expenditure, their analyses employed the cross-sectional variations in the pooled cross-sectional sample. An exceptional study is Saaritsa (2011), which shows the estimated elasticity to temporary income from a fixed effect one-way error component model using a longitudinal household budget survey in 1928 Helsinki. Hence, the estimated elasticity tells the mean tendency that the households with lower (higher) income were more (less) likely to use a specific risk-coping strategy. [^8] In other words, the elasticity is derived from the observations across different households with different income levels. Since risk-coping is a concept about the dynamic behavior of a specific household, however, comparing the different time cells would be more natural to evaluate the roles of the risk-coping strategies. By using the panel dataset, this study implements the stylized econometric method to systematically investigate the households’ smoothing mechanisms to offer comparable estimates.
Another historical contribution is adding micro-level evidence on consumer behaviors during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^9] There are several business history studies analyzing the relationship between consumption and production. A representative work is Gordon (2013), who investigated the expansion of the sewing machine from many aspects, including the production, distribution, and gender social norms in Japan. However, these studies have little interest in the households’ risk-coping behaviors. A few studies investigated on the relationships between household chores and consumption behaviors in interwar households (Tanimoto 2011, 2012). This study employs a different analytical view by focusing on short-run consumption behaviors. Looking into the extent and how the local society could ensure the households’ consumption would provide valuable insights into the consumption history literature.
The rest of this paper is organized as follows. Section 2 overviews the Japanese economy around the First World War (WWI) and reviews Tsukisima, the target area in my analysis. I also summarized the financial institutions that were accessible to working-class households. Section 3 introduces the documents used to construct my panel dataset and discusses the sample characteristics. Section 4 summarizes the empirical strategies and results. Section 5 discusses the roles of credit purchases in the local economy. Section 6 concludes.
@@ 2 Historical Background
2.1 Japanese Economy around the First World War
Japan had the foundations of an industrial nation by WWI. Between 1885 and 1915, real GNP grew annually by 2.6%, while population increased at 1.1%. [^10] Given that economic growth is faster than population growth, this implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively.
@@ 2.1 Japanese Economy around the First World War
Soon after the outbreak of WWI, exports from European countries declined. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus. Given the declines in the export capacity of the European countries, import substitution had progressed. This expands the domestic market and leads to a rise in the machinery, metal, and chemical sectors. A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated as a leading industry. The development of the shipbuilding industry stimulated demand for the machinery and steel industries. Through this process of heavy and chemical industrialization, the total number of enterprises increased. [^11] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p. 128), Nakamura and Odaka (1989, p. 10), and Sawai and Tanimoto (2016, p. 250). Year-on-year growth in the number of incorporated firms was 17.8% in 1918, 22.3% in 1919, and 13.5% in 1920. These are the top three years of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). Figure 1 shows that the GNP growth from 1915-19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
@@
This rapid industrialization impacted the Japanese economy in many ways. Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar. Although wages began to rise around 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished. Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities. The share of large cities (with people of 50,000 or more) as a percentage of the total population was less than 10% in 1889 but increased to nearly 16% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during the WWI period, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems. [^12] Descriptions on the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp. 21–23) and Nakamura (1971, pp. 142; 147).
@@
Exports in the 1920s were lower than before because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign species during WWI. Since the current account deficit was settled through the disbursement of foreign species owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices, especially prices of nontradable goods, being kept high. The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. [^13] On the other hand, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p. 145). Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp. 253–254).
@@
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934-1935 price.
Sources: Created by the author using Ohkawa et. al. (1974, p. 237).
As described, there were short-term business cycles, such as the boom caused by WWI and the subsequent recession. However, taken as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century and before WWII. As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915-19 and 118% in 1920-24.141414Total consumption as a percentage of the GNP remained above 80%. Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never turned negative except during the Russo-Japanese War period. This was a high level even by international standards. Given the rapid urbanization, therefore, it is plausible to focus on the consumption behavior of urban households when considering the Japanese economy in this period. However, progress has yet to be made. [^15] Descriptions on the personal consumpsion expenditure are from Nakamura (1971, pp. 2–5; 27; 30–31).

### v10

# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

@@ Abstract
I analyzed the risk-coping strategies among factory worker households in early 20th-century Tokyo. I digitized and analyzed a unique daily longitudinal household budget survey to determine how consumption was impacted by idiosyncratic shocks. I found that while the households were so vulnerable that the shocks impacted their consumption levels, the estimated income elasticity for food consumption was relatively low in the short run. The event-study analysis using adverse health shock confirms the robustness of the results. The result of mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. Despite the potential loss of profit, retailers in a competitive situation allowed consumers to trade on credit. This shows the roles of informal credit institutions in mitigating vulnerability among urban worker households.
Keywords: Consumption smoothing; credit institution; urban local community; measurement error; risk-coping strategy; risk-sharing
@@ 1 Introduction
Economic theory illustrates the mechanism behind the smoothing of consumption in complete markets (Arrow and Debreu 1954). Additionally, a series of empirical studies have tested the smoothness of consumption among households in various economies under systematic empirical specification. Evidence shows that markets in contemporary developing countries are often imperfect.111 The early representative literature includes, for example, Townsend (1994), Townsend (1995), and Ravallion and Chaudhuri (1997). Morduch (1999), Townsend (1995), and Dercon (2004) provide reviews of the earlier studies. Attanasio and Pistaferri (2016) and Meyer and Sullivan (2023) provide recent discussions. The households are vulnerable to risk, and idiosyncratic shocks impact the levels of their consumption because shocks are not fully insured in the markets. Therefore, consumption smoothing is an essential mechanism for mitigating poverty, especially in developing economies. [^2] Smoothness of consumption is also an important factor in human capital formation. For example, Foster (1995) investigates the shock of the 1988 floods in Bangladesh and shows that children’s growth patterns are influenced by the inefficiency of the credit markets. Rose (1999) investigates the relationship between consumption smoothing and excess losses of girls, revealing that the favorable rainfall events in early life increased the survival probabilities of girls in rural India. Gertler and Gruber (2002) also shows that consumption smoothing against health shocks is imperfect in Indonesia.
Credit institutions in local society can mitigate idiosyncratic income shocks, especially when the availability of formal financial institutions is limited. Udry (1994) showed that state-contingent loans among relatives and neighbors could partially insure against idiosyncratic shocks on rural households in northern Nigeria. A detailed field survey revealed that asymmetric information was no longer important in those private credit contracts (Udry 1990). Fafchamps and Lund (2003) also presented evidence that the households in the rural Philippines used informal networks with relatives and friends as a risk-mitigation strategy for idiosyncratic shocks. Related studies consider risk-sharing among households with similar cultural and institutional backgrounds. Grimard (1997) showed that insurance among households in the same ethnic group could partially mitigate idiosyncratic shocks in the multiethnic economy of Côte d’Ivoire. Munshi and Rosenzweig (2016) presented the importance of rural insurance networks among workers in the same caste in the context of spatial wage inequality in India. The evidence suggests that the availability of informal credit institutions relates to the stability of consumption in rural economies.
This study assesses risk-coping behavior in informal credit transactions in the urban community in early twentieth-century Tokyo. I digitized and analyzed a unique daily longitudinal budget survey on factory worker households conducted in a representative manufacturing area. I found that the full insurance hypothesis is strongly rejected for overall consumption expenditure, meaning that the factory worker households in the historical metropolis suffered certain consumption losses when they faced idiosyncratic adverse income shocks. However, expenses on indispensable consumption categories were partially smoothed in the short run. Credit purchases in local retailers played an important role in mitigating adverse idiosyncratic shocks, whereas the other risk-mitigation strategies were impractical. Credit purchases did not require an explicit contract and thus often led to a loss in sales for the retailers. However, to remain competitive, retailers allowed credit transactions for their customers to remain regular customers. This shows that informal credit institutions in urban local communities functioned as informal insurance networks in rural economies.
@@
The first contribution of this study is to offer a case study of a metropolis in a developing economy. Urbanization is now a worldwide phenomenon. Moreover, the proportions of urban population in many developing countries, especially in East Asia, are reaching a similar level to the developed countries (Henderson and Turner 2020). This requires a deeper understanding of consumption behavior among urban working households in the economic development process. However, risk-mitigation strategies among urban worker households have been understudied. [^3] Extant studies have focused particularly on the rural economies (e.g., Townsend 1994; Ravallion and Chaudhuri 1997; Gertler and Gruber 2002). Recent field experimental studies also focus on village economies (Comola and Prina 2023; Macours et al. 2022). Although several studies considered consumption smoothing among urban households in postwar Bangkok and prewar Osaka, the risk-mitigation mechanisms, particularly via credit institutions, are not fully evaluated. [^4] See Townsend (1995) and Ogasawara (2024), respectively. Osaka was middle scale city of approximately half the size of Tokyo City circa 1920 (Online Appendix Table A.1). (Skoufias and Quisumbing 2005; Aguila et al. 2017) also provided case studies on middle-sized contemporary economies. However, they did not consider the roles of local credit institutions. This study considers a set of homogeneous households in a representative manufacturing area in prewar Tokyo, which provides a valuable empirical setting for testing risk-sharing strategies in the historical metropolis. [^5] Focusing on a sample that shares similar occupational and regional characteristics systematically mitigates bias in estimating income elasticity under the standard risk-sharing regressions due to the heterogeneity in the risk preferences (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). My findings show that informal credit institutions in the local urban community could mitigate vulnerability among households similar to rural developing economies (Udry 1994; Fafchamps and Lund 2003).
@@
The second contribution is to add analytical perspectives to economic history studies. First, studies of how working-class households overcame their economic hardships during the late 19th to the early 20th century have focused primarily on Britain. The weight of evidence supports the roles of historical financial institutions and labor supply adjustments in compensating for income loss in working-class households. [^6] Horrell and Oxley (2000) highlight the important role of labor supply adjustments among family members to manage the income losses of the head in working-class households in late 19th-century Britain. Scott and Walker (2012) also found that interwar British working-class households had access to various institutions such as clothing clubs, hire-purchase, contractual savings, and insurance. The households with less earnings were likely to use credit-based institutions like clubs and hire-purchase, whereas those with higher incomes could use savings institutions. Kiesling (1996) and Boyer (1997) focused on the risk-mitigation strategies among the cotton workers households in Lancashire during the recession caused by the Civil War. However, there is little evidence of how those risk-coping strategies smoothed their consumption. Similarly, previous studies have used cross-sectional variations in income to estimate the elasticities of the temporary income sources. [^7] For example, Horrell and Oxley (2000) and James and Suto (2011) consider the deviations of actual earnings from the predicted earnings as the shocks. Scott and Walker (2012) employed the cross-sectional variations in the pooled cross-sectional sample to analyze the smoothness of expenditure (deviation from mean expenditure). An exceptional study is Saaritsa (2011), which shows the estimated elasticity of temporary income from a fixed effect one-way error component model using a panel household budget survey in 1928 Helsinki. The estimated elasticity is, therefore, potentially disturbed by household-specific unobservable confounding factors such as preference and learning ability. As risk-mitigation is a concept in the dynamic behavior of a specific household, the estimator derived from within variations of the units is more suitable for evaluating the roles of risk-mitigation strategies. This study implements a stylized econometric method for the panel dataset to offer comparable estimates for future studies. Second, this study adds micro-level evidence on consumer strategies during industrialization. Francks (2009) provides a comprehensive view of the long-run consumption history in Japan from the Edo era to the twentieth century. [^8] Several business history studies analyze the relationship between consumption and production. A representative work is Gordon (2013), which investigated the expansion of the sewing machine from many perspectives, including production, distribution, and gender social norms in Japan. However, these studies have not focused on the households’ risk-mitigation strategies. Some studies investigated the relationships between household chores and consumption strategies in interwar households (Tanimoto 2011, 2012). This study employs an alternative analytical view by focusing on short-run consumption strategies. An examination of how, and the extent to which, the local society could ensure household consumption provides valuable insights into consumption history literature. Third, several financial historical studies provide supply-side (lender-side) evidence on the functions of the micro-financial institutions for the workers (O’Connell and Reid 2005). This study is the first to document the function of informal credit transactions by local retailers in the context of consumption smoothing. Munshi and Rosenzweig (2016) linked spatial wage inequality and men’s mobility in India with the availability of risk-mitigation networks among workers in the same caste. From this perspective, lower barriers of local credit institutions may motivate rural workers to urbanize, which may partially explain why Tokyo grew so fast circa WWI. This hypothesis requires testing in future research.
@@ 2 Historical Background
2.1 Japanese Economy circa the First World War
@@
Between 1885 and 1915, the Japanese real GNP grew annually by 2.6%, whereas the population increased by 1.1%. [^9] Faster economic than population growth implies that modern economic growth had started in this period as defined by Simon Kuznets (Miyamoto 2008, pp. 60; 64). In 1915, the percentage of the employed population in the primary, secondary, and tertiary industries were 63%, 20%, and 18%, respectively. Based on domestic genuine production, these industries accounted for just under 30%, over 30%, and approximately 40%, respectively. Soon after the outbreak of WWI, exports from European countries declined. Japan increased exports to the Asian market on behalf of those European countries and exported munitions and foodstuffs to the European Allies, leading to a current account surplus. The import substitution expanded the domestic market and led to a rise in the machinery, metal, and chemical sectors. A worldwide shortage of ships also stimulated the development of shipbuilding, which proliferated. The development of the shipbuilding industry further stimulated demand for the machinery and steel industries. Through this process of heavy and chemical industrialization, the total number of enterprises increased. [^10] Descriptions about the process of heavy chemical industrialization are based on Nakamura (1971, p. 128), Nakamura and Odaka (1989, p. 10), and Sawai and Tanimoto (2016, p. 250). Year-on-year growth in the number of incorporated firms was 17.8%, 22.3%, and 13.5% in 1918, 1919, and 1920, respectively. These are the top three of the 34 years from after the Russo-Japanese War to 1940 (Nakamura and Odaka 1989, p. 22). Figure 1 shows that the GNP growth from 1915–19 was 11.5% in nominal terms (4.4% in real terms), the highest growth throughout the 1910s and 1920s.
@@
This rapid industrialization impacted the Japanese economy in many ways. Due to export-led economic growth, large inflows of gold increased domestic currency balances, causing prices to soar. Although wages began to rise circa 1916 as a response to rising prices and labor demand due to the increased number of firms, wage increases failed to keep pace with soaring prices. By the end of the war, the number of labor disputes grew sharply, and the labor movement flourished. Furthermore, the rise of business accelerated urbanization as the rural population moved to cities in search of job opportunities. The share of large cities (with a population of $50,000$ or more) as a percentage of the total population was less than $10$% in 1889 but increased to nearly $16$% in 1920 (Online Appendix Figure A.1a). The growth rate was exceptionally high during WWI, indicating that structural changes through the war accelerated the rise of the urban population. Urbanization has increased public investment in large cities, especially in construction investment in social capital such as roads and water supply systems. [^11] Descriptions of the export-led growths, urbanization, and investments are from Nakamura and Odaka (1989, pp. 21–23) and Nakamura (1971, pp. 142; 147).
@@
Exports in the 1920s were lower because of the appreciation of the yen against the real exchange rate due to the accumulation of domestic and foreign currencies during WWI. As the current account deficit was settled through the disbursement of foreign currencies owned by the government and the Bank of Japan, there was no reduction in the domestic money supply despite the sizeable current account deficit. This resulted in the domestic prices (especially prices of non-tradable goods) remaining high. The first half of the 1920s saw a temporary decline in industrial investment as the capital investment boom of the WWI period came to a halt under declining demand. [^12] Conversely, construction investment by the private sector, including electric power and private railways, increased (Nakamura 1971, p. 145). Descriptions about the post-war economy in this paragraph are from Sawai and Tanimoto (2016, pp. 253–254).
@@
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934–1935 price.
Source: Created by the author using Ohkawa et. al. (1974, p. 237).
As described, there was a short-term business cycle characterized by the boom caused by WWI and the subsequent recession. However, as a whole, Japan achieved a high economic growth rate relative to other countries in the early 20th century. As shown in Figure 1, the growth rates of personal consumption expenditure were particularly high during and immediately after WWI, with average growth rates reaching 112% in 1915–19 and 118% in 1920–24. This was a high level by international standards. [^13] Total consumption as a percentage of the GNP remained above 80%. Although this level is not apparent among other countries (Kuznets 1962), the growth rate of real personal consumption per capita was relatively high: from 1875 to 1939, the average growth rate of personal consumption was 1.36%, and the rates never became negative except during the Russo-Japanese War period. Descriptions on the personal consumption expenditure are from Nakamura (1971, pp. 2–5; 27; 30–31). This highlights the importance of revealing the consumption behavior of urban households to understand the mechanisms underpinning the rapid Japanese economic growth in this period.